OpenStudy (anonymous):

Solve For Y ; 6x+2y=20 . Right Now I Got ; 2y=20-6x . I Keep Getting Suck Right Here

OpenStudy (anonymous):

*Stuck

OpenStudy (cgreenwade2000):

Divide both sides of the equation by 2. That will give you y=.......

OpenStudy (anonymous):

\[6x+2y=20\\3x+y=10\\y=10-3x\]

OpenStudy (anonymous):

y=10-3x ?

OpenStudy (anonymous):

Yeaah That What I Got .

OpenStudy (cgreenwade2000):

@oldrin.bataku Why give the answer? There was already a method given to solve the problem.

OpenStudy (anonymous):

@cgreenwade2000 What do I Do Once I Get To ; y=10-3x ?

OpenStudy (anonymous):

@cgreenwade2000 I showed how to do it by dividing by 2 first, which is simpler in this case.

OpenStudy (cgreenwade2000):

@Isabel.A by getting y=...... You have solved for y. You now know what y is equal to. @oldrin.bataku That's fine. But make sure you help out and not give answers. That takes away from them learning the steps.

OpenStudy (anonymous):

No ._. I;m Stuck At y=10-3x .

OpenStudy (cgreenwade2000):

The question says to solve for Y. Y=10-3X is already solved for Y.

OpenStudy (cgreenwade2000):

If I ask you to tell me what Y is equal to, then you would say it equals 10-3x.

OpenStudy (anonymous):

Ohhh . Okay so That As Far As It Can Go ?

OpenStudy (cgreenwade2000):

You can not solve anything else for y. There are no steps left.

OpenStudy (anonymous):

Okay(: Thank You Sooooooo Much For The Help Think I Got It Now .

OpenStudy (anonymous):

@oldrin.bataku Thank You (: